lim sin-1 2x/x
x-->0

lim (csc x- 1/x)
x-->0

show step by step
1
Expert's answer
2020-09-19T13:00:19-0400

For first pat

"\lim_{x \to0} \frac{Sin^{-1}2x}{x}"


Apply L.hospital rule in above equation,

"\lim_{x \to 0} \frac{2}{\sqrt{1-4x^2}}"

Now putting the limit

we get

"\frac{2}{\sqrt{1-0}}"

=2


Now second part the given limit is

"\lim_{x \to0} \frac{cosec x^{-1}}{x}"


"\lim_{x \to0} \frac{cosec \frac{1}{x}}{x}"

This limit does not hold good for L.hospital rule, So the given limit diverges to "\infty"